July 23-28, 2007 LGS observing run

Night logs and results

7/24/07 UT
Clear, excellent seeing. Completed laser and BTO setup for first 4 hrs. Then measured bright star performance (~30% Strehl at K), observed several science targets, and measured vignetting on LOWFS. Laser stable at 5.0 W, FWHM=2.3". palm_lgs_log_070724.pdf

7/26/07 UT
Partly cloudy, excellent seeing. AO acquisition camera power supply failed. Used temporary power supply and software integration of video signal. Unable to focus LLT or acquire faint stars. Observed one LGS and several NGS science targets. Laser 5.0W at start, 4.5 W and end of night. palm_lgs_log_070726.pdf

7/27/07 UT
Partly cloudy, excellent seeing. 6 AO software crashes during night. Observed 2 LGS targets and several NGS backup targets, with frequent interruption by clouds. Laser power 7.0 W at start, 6.9 W at end of night, FWHM=2.6". LGS return significantly improved (200 cts @ 150 Hz.) palm_lgs_log_070727.pdf

7/28/07 UT
Mostly cloudy, good seeing. No AO software crashes; used IDL background routine. Observed 4 LGS targets through thin clouds, with frequent interruptions by clouds (lost ~4hr). Laser power 7.2 W at start, 7.0 W at end of night, FWHM=2.4". Laser AOM driver failed at 12:35am, but Renu brought it back to life. LGS performance good despite clouds. palm_lgs_log_070728.pdf

7/29/07 UT
Partly cloudy, excellent seeing. No AO software crashes. Observed 5 LGS targets, with several interruptions by clouds (lost ~2hr). Laser power 7.0 W at start, 6.9 W at end of night, FWHM=2.2". Laser AOM driver failed again at 03:15am, but Renu brought it back to life. Set new LGS Strehl record of ~45% in K (to be confirmed). Excellent end to a cloudy observing run. palm_lgs_log_070729.pdf

ASCAM movies

data
missing
7/24/07 UT 7/26/07 UT 7/27/07 UT 7/28/07 UT 7/29/07 UT

Safety documents

Personnel and engineering schedules

  Mon. 7/23 Tue. 7/24
(NGS)
Wed. 7/25 Thu. 7/26 Fri. 7/27 Sat. 7/28
Angione X X X X X X
Bouchez X X X X X X
Cromer X X X X X X
Guiwits X          
Martin   X X X X X
Petrie X X        
Roberts X X X X ?  
Tripathi X X X X X X
Trinh X          
Troy         X X

  • Engineering Schedule:

Procedures and test plans

Please post your test plans on the Test Plan Library page, and add a link below.
  • AO setup
    • AO system internal focus: doc, pdf
    • AO bench setup and checkout: doc, pdf
  • BTO setup
    • BTO setup (v2.3): doc, pdf
    • BTO checkout: doc (v2.3), pdf (v2.3)
    • BTO shutdown (v1.1): doc, pdf
  • Observing procedures:
    • NGS Checkout (v3.1): doc, pdf
    • Laser Acquisition (v5.1): doc, pdf
    • LGS target acquisition (v3.5): doc, pdf
    • LGS target acquisition (v4.0): doc, pdf
  • Engineering experiments:
    • AO Optimization on the laser at zenith: doc, pdf
    • Low-order loop optimization: doc, pdf
Topic attachments
I Attachment History Action Size Date Who Comment
Texttxt PMR-palomar070724.txt r1 manage 16.1 K 2007-07-20 - 18:37 AntoninBouchez  
Texttxt PMR-palomar070726.txt r1 manage 16.1 K 2007-07-20 - 18:37 AntoninBouchez  
PDFpdf lgs070723_schedule_v1-0.pdf r1 manage 63.3 K 2007-07-18 - 18:02 AntoninBouchez  
Microsoft Excel Spreadsheetxls lgs070723_schedule_v1-0.xls r1 manage 531.5 K 2007-07-18 - 18:02 AntoninBouchez  
PDFpdf palm_lgs_log_070724.pdf r1 manage 53.9 K 2007-07-26 - 21:40 AntoninBouchez  
PDFpdf palm_lgs_log_070727.pdf r2 r1 manage 47.4 K 2007-07-27 - 21:51 AntoninBouchez  
PDFpdf palm_lgs_log_070728.pdf r1 manage 45.5 K 2007-07-29 - 03:09 AntoninBouchez  
PDFpdf palm_lgs_log_070729.pdf r1 manage 49.9 K 2007-07-29 - 12:04 AntoninBouchez  
Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | Raw edit | More topic actions
Topic revision: r17 - 2007-07-31 - AntoninBouchez
 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback